Opened 10 years ago

Closed 9 years ago

Last modified 3 years ago

#263 closed flight processing (fixed)

IPY09/11, flight day 219/2009, Helheim

Reported by: knpa Owned by: knpa
Priority: alpha 4 high Milestone:
Component: Archiving Keywords:
Cc: Other processors:

Description (last modified by mark1)

Data location: ~arsf/arsf_data/2009/flight_data/ipy/IPY09_11-2009_219_Helheim

Data arrived from ARSF via SATA disk Daphne03 on 24/08/09.

Scientific objective: Ice thickness measurements to add to a record of two major glaciers and to create a high res DEM.

Alpha 4-High

PI: T. Murray

Sensors:

  • Eagle (08/04/2010)
  • Leica LIDAR (08/04/2010)
  • Digital Photography (11/02/2010)

Change History (19)

comment:1 Changed 10 years ago by knpa

  • Description modified (diff)

comment:2 Changed 10 years ago by knpa

Corrected logsheet project code from GB09_21

comment:3 Changed 10 years ago by benj

For day 215 I got much better IPAS results by downloading data from the hel2 station than by using Kulu. I suggest whoever processes the IPAS (+Applanix) for this does the same.

comment:4 Changed 10 years ago by mark1

  • Description modified (diff)

Photographs dispatched to PI on 11/02/2010

comment:5 Changed 10 years ago by mark1

Started processing the LiDAR and Hyperspectral

comment:6 Changed 10 years ago by mark1

  • Owner set to mark1
  • Status changed from new to assigned

comment:7 Changed 10 years ago by emca

Delivery checked LIDAR data at
~airborne/workspace/IPY09_11-2009_219_Helheim/delivery/20100325
All looks good. Ready to deliver.

comment:8 Changed 10 years ago by emca

Delivery checked Hyperspectral data at
~airborne/workspace/IPY09_11-2009_219_Helheim/delivery/20100326
and all is good. Ready to deliver.

comment:9 Changed 10 years ago by mark1

rsynced data to main ~arsf repository.

comment:10 Changed 10 years ago by mark1

SCT offsets for Eagle data:
line 1: 0.0
line 2: 0.04

comment:11 Changed 10 years ago by mark1

  • Description modified (diff)

Hyperspectral and LIDAR data delivered to PI on (08/04/2010)

comment:12 Changed 9 years ago by mark1

  • Component changed from Processing: general to Archiving

All data delivered, ready for archiving.

comment:13 Changed 9 years ago by knpa

  • Owner changed from mark1 to knpa

Started archiving.

comment:14 Changed 9 years ago by knpa

There are some missing applanix files, specifically the contents of gps (POSGNSS associated files).

comment:15 Changed 9 years ago by knpa

raw applanix files (DCALM), start at 29 for some reason?

comment:16 Changed 9 years ago by knpa

Finished archiving prep, tarball ready to be downloaded by NEODC.

comment:17 Changed 9 years ago by knpa

Backed up by NEODC, transferred to archive disk.

comment:18 Changed 9 years ago by knpa

  • Resolution set to fixed
  • Status changed from assigned to closed

Deleted workspace copy.

Closing ticket.

comment:19 Changed 3 years ago by asm

Archiving

Re-processed hyperspectral data has been archived and everythin is ready to be downloaded from CEDA. Will close this ticket.

Note: See TracTickets for help on using tickets.